3x+1=6. how do i find out what x is?

Answer by sdmmadam@yahoo.com(530)

3x+1=6. how do i find out what x is?
3x+1=6. ----(1)
This is a linear equation in x.
(that is an equation in the variable x with the highest power for x equal to 1)
Here x is the unknown and three times x added to 1 is 6
(3x+1)-1 = 6-1 (Subtracting 1 from both the sides)
3x+(1-1) = 5 (by additive associativity)
3x+0 =5
3x= 5
x= 5/3
Answer: x = 5/3
Verification: Put x = 5/3 in (1)
LHS = 3x+1 = 3X(5/3)+1 = 5+1 = 6 = RHS
Therefore our value for x is correct.
Note: why did we subtract 1 both the sides?
We observe 1 added to 3x on the left side and inorder to retain only the term containing the unknown x we have to remove mathematically the term 1 from the left side. So we add (-1)to both the sides or subtract 1 from both the sides toget rid of 1 from the left.
Though we have so much mathematics behind the whole thing,we may practically handle the problem the following way
Consider 3x+1=6.
Taking 1 from the left side to the right side following the rule : change side then change sign
3x= 6-1
3x=5
three multiplied by x is 5
Therefore divide both sides to get (1)x that is our unknown x.
Got it!